OpenStudy (henrietepurina):

Solve the system of equations by substitution. What is the solution for x? 2x + y = 1 4x + 2y = -1

OpenStudy (henrietepurina):

x = 0 x = 2 There is no x value as there is no solution. x can be any value as there is an infinite number of solutions.

OpenStudy (henrietepurina):

its like a b c d

OpenStudy (anonymous):

ok let's solve the first equation for y: 2x + y = 1 y = 1 - 2x

OpenStudy (anonymous):

Then we substitute this into the other equation: 4x + 2(1-2x) =-1

OpenStudy (anonymous):

Can you solve it from here?

OpenStudy (henrietepurina):

4x + 2 - 4x = -1?

OpenStudy (henrietepurina):

4x - 4x = 1?

OpenStudy (henrietepurina):

0 = 1?

OpenStudy (anonymous):

Yup, and what does that mean?

OpenStudy (henrietepurina):

there is no answer or in other words c?

OpenStudy (anonymous):

Nicely done :)

OpenStudy (henrietepurina):

Can you help with another?

OpenStudy (anonymous):

ok

OpenStudy (henrietepurina):

Solve the system of equations. 5x + y = 9 3x + 2y = 4

OpenStudy (henrietepurina):

(-2, 5) (1, 4) (2, -1) (4, -4)

OpenStudy (henrietepurina):

let me try I think I can do it and then you check!

OpenStudy (henrietepurina):

is it c?

OpenStudy (anonymous):

Yes it is

OpenStudy (henrietepurina):

thank you :) howd you get it?